# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ 4 Gaussian Width of Deep Learning Gradients
In this section, we initiate the study of loss gradient Gaussian width (LGGW) of deep learning gradients. As a warm up, we review that bounded $L_{2}$ norm, e.g., $\|\xi(\theta;z)\|_{2}\leq c$ for any $\theta\in\Theta$ and $z\in{\cal Z}$, with no additional structure, leads to a large Gaussian width since for such $\hat{\Xi}_{n}$, $w({\hat{\Xi}_{n}})=\mathbb{E}_{\mathbf{g}}[\sup_{\hat{\xi}_{n}}\langle\hat{\xi}_{n},\mathbf{g}\rangle]\leq c\mathbb{E}_{\mathbf{g}}[\langle\mathbf{g}/\|\mathbf{g}\|_{2},\mathbf{g}\rangle]=c\mathbb{E}_{\mathbf{g}}[\|\mathbf{g}\|_{2}]=O(\sqrt{p})$. We posit that the Gaussian width of deep learning gradients is much smaller and present a set of technical results in support. In Section 4.1, motivated by the empirically observed sharp eigenvalue decay of loss Hessians, under mild assumptions we show that LGGW can be bounded by an upper bound on the $L_{2}$ norm of the Hessian eigenvalues, which has been empirically shown to be $\tilde{O}(1)$ [92, 87]. Further, in Section 4.2, we show that under standard assumptions, the LGGW can be bounded in terms of the Gaussian width of the featurizer, i.e., output of the last-but-one layer, for both feedforward and residual networks. The results emphasize the benefits of Hessian eigenvalue decay and the stability of featurizers for deep learning.
4.1 Loss Gradient Gaussian Width using Empirical Eigen Geometry of Hessian
In the context of deep learning, recent work has illustrated that empirically the gradients are considerably structured and gradients are effectively live in a low-dimensional space [33, 64, 65, 29, 47], and close to convergence, almost all gradients are really close to zero, the so-called interpolation condition [13, 62, 55, 10]. Such observations have formed the basis of more memory or communication-efficient algorithms, e.g., based on top-$k$ count-sketch, truncated SGD, and heavy hitters based analysis [71, 39]. In parallel, considerable advances have been made in understanding the eigenspectrum of the Hessian [92, 87, 89, 48]. In particular, fast decaying eigenvalues have been observed across the entire range of models including ConvNets, ResNets, as well as GPT and ViT variants [92, 89], and some work [87] have empirically shown that the $k$-th largest (absolute) eigenvalue decays following power-law, i.e., $\lambda_{k}\leq\lambda_{1}k^{-s}$ with $s>\frac{1}{2}$ empirically.
We make some assumptions based on the empirically observed eigenvalue decay of the Hessian and establish bounds on the loss gradient Gaussian width based on properties of the Hessian eigenvalues.
Assumption 3.
Let $\hat{H}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\ell(\theta,z_{i})$ denote the loss Hessian for $\theta\in\Theta$ and let $\bm{\lambda}(\theta)=[\lambda_{i}(\theta)],i\in[p]$ denote the (not necessarily positive) eigenvalues of $\hat{H}_{n}(\theta)$. For any $\theta\in\Theta$, the empirical gradient $\hat{\xi}_{n}(\theta)=\hat{H}_{n}(\tilde{\theta})(\theta-\theta_{0})+\hat{\xi}_{n}(\theta_{0})$ for some $\theta_{0}\in\mathbb{R}^{p},\|\theta-\theta_{0}\|_{2}\leq c$, where $\tilde{\theta}=a_{\theta}\theta+(1-a_{\theta})\theta_{0},a_{\theta}\in[0,1]$. With $\bm{\lambda}(\tilde{\theta})$ denoting the Hessian eigenvalues of $\hat{H}_{n}(\tilde{\theta})$ at any such $\tilde{\theta}$, we assume $\left\|\bm{\lambda}(\tilde{\theta})\right\|_{2}\leq\Lambda^{(u)}$ for some fixed $\Lambda^{(u)}\in\mathbb{R}_{+}$.
Remark 4.1.
Assumption 3 is inspired by the empirical evidence of Hessian eigenvalue decay especially close to convergence [92, 87], and it is reasonable to expect $\left\|\lambda(\tilde{\theta})\right\|_{2}=\tilde{O}(1)$, which means we could set $\Lambda^{(u)}=\tilde{O}(1)$. In Assumption 3, $\theta_{0}$ can be from a pretrained model or a random initialization. Further, we can replace the single $\theta_{0}$ by a set $\Theta_{0}$ of cardinality $|\Theta_{0}|=N$, so that for any $\theta\in\Theta,\exists\theta_{0}\in\Theta_{0}$ such that $\|\theta-\theta_{0}\|_{2}\leq c$, and our main result below will continue to hold with only an additional $\log N$ term [58, 17]. ∎
Theorem 4.
Under Assumption 3, the LGGW for $\theta\in\Theta$ is $w(\hat{\Xi}_{n})\leq c\Lambda^{(u)}\sqrt{\log p}$.
Remark 4.2.
The width bound in Theorem 4 depend on $\Lambda^{(u)}$, the upper bound of $\left\|\bm{\lambda}(\tilde{\theta})\right\|_{2}$, so that a power law decay of sorted absolute Hessian eigenvalues $|\lambda_{(i)}(\tilde{\theta})|\leq c^{(u)}i^{-s}$ with $s\geq\frac{1}{2}$ (e.g., as empirically observed in [87]) will lead to $\Lambda^{(u)}=O(\sqrt{\log p})$. Then, the overall bound is $w(\hat{\Xi}_{n})\leq c\log p$ yielding a generalization bound from (11) with complexity term $c\frac{\log p}{\sqrt{n}}$, hence a sharp bound. ∎
@@ 4.2 Gaussian Width Bounds for Neural Networks
Let $\theta_{0}:=(\operatorname{vec}(W_{0}^{(1)})^{\top},\ldots,\operatorname{vec}(W_{0}^{(L)})^{\top},\mathbf{v}_{0}^{\top})^{\top}$ denote the initial weights, and $w_{0,ij}^{(l)}$ denote $i,j$-th element of weight matrix $W_{0}^{(l)}$ for $l\in[L]$. At initialization, $w_{0,ij}^{(l)}\sim{\cal N}(0,\sigma_{0}^{2})$ for $l\in[L]$ where $\sigma_{0}=\frac{\sigma_{1}}{2\left(1+\frac{\sqrt{\log m}}{\sqrt{2m}}\right)},\sigma_{1}>0$, and $\left\lVert\mathbf{v}_{0}\right\rVert_{2}=1$.
@@
Assumption 1 is satisfied by commonly used activation functions such as (non-smooth) ReLU [2] and (smooth) GeLU [38]. The second assumption is regarding the random initialization of the weights, variants of which are typically used in most work in deep learning theory [40, 2, 37, 30, 79, 6, 22, 21].
4.2.1 Bounds for FeedForward Networks (FFNs)
We consider $f$ to be an FFN given by
$f(\theta;x)=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m}}W^{(1)}\mathbf{x}))))~{},$ (22)
where $W^{(1)}\in\mathbb{R}^{m\times d},W^{(l)}\in\mathbb{R}^{m\times m},l\in\{2,\ldots,L\}$ are the layer-wise weight matrices, $\mathbf{v}\in\mathbb{R}^{m}$ is the last layer vector, $\phi(\cdot)$ is the (pointwise) activation function, and the total set of parameters $\theta\in\mathbb{R}^{p},p=(L-1)m^{2}+md+m$ is given by
Our analysis for both FFNs and ResNets will be for gradients over all parameters in a fixed radius spectral norm ball around the initialization $\theta_{0}$, which is more flexible than Frobenius norm balls typically used in the literature [85, 19, 94, 8]:
[…]
Let $B_{\rho}^{\operatorname{Spec}}(\theta_{0})\in\mathbb{R}^{p-m}$ be the corresponding set of weight matrices $W=\{W^{(l)},\ell\in[L]\}$. The main implications of our results and analyses go through as long as the layerwise spectral radius $\rho=o(\sqrt{m})$ and last layer radius $\rho_{1}=O(1)$, which are arguably practical.
Let
$h^{(L)}(W,\mathbf{x})$ $:=\phi\left(\frac{1}{\sqrt{m}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m}}W^{(1)}\mathbf{x})))\right)~{}\in\mathbb{R}^{m}$ (25)
be the “featurizer,” i.e., output of the last but one layer, so that $f(\theta;\mathbf{x})=\mathbf{v}^{\top}h^{(L)}(W,\mathbf{x})$ with $\theta=(W,\mathbf{v})$. Consider the set of all featurizer values for a single sample and as average over samples:
$A^{(L)}(\mathbf{x})=\left\{h^{(L)}(W,\mathbf{x})\mid W\in B_{\rho}^{\operatorname{Spec}}(\theta_{0})\right\},~{}\hat{A}_{n}^{(L)}:=\left\{\frac{1}{n}\sum_{i=1}^{n}h^{(L)}(W,\mathbf{x}_{i})\mid W\in B_{\rho}^{\operatorname{Spec}}(\theta_{0})\right\}~{}.$ (26)
We now establish an upper bound on the LGGW of FFN gradient set $\hat{\Xi}_{n}^{\operatorname{ffn}}$ corresponding to parameters in $B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$. In particular, we show that the LGGW can effectively be bounded by the Gaussian width of the sample average of the featurizer.
Theorem 5 (Gaussian Width of Gradients: FFNs).
Under Assumptions 4 and 5, with $\beta:=\sigma_{1}+\frac{\rho}{\sqrt{m}}$, with probability at least $\left(1-\frac{2L}{m}\right)$ over the randomness of the initialization, we have
$w(\hat{\Xi}_{n}^{\operatorname{ffn}})\leq c_{1}w(\hat{A}_{n}^{(L)})+c_{2}L(1+\rho_{1})\beta^{L-1}~{}.$ (27)
Remark 4.3.
Note that choosing $\sigma_{1}<1-\frac{\rho}{\sqrt{m}}$, i.e., mildly small initialization variance, satisfies $\beta<1$. When $\beta<1$, we have $w(\hat{\Xi}_{n}^{\operatorname{ffn}})=O(1)$ as long as $L=\Omega(\frac{\log(\sqrt{m}+L(1+\rho_{1}))}{\log\frac{1}{\beta}})$, which holds for moderate depth $L$. To see this, note that we have $w(\hat{\Xi}_{n}^{\operatorname{ffn}})=O(w(\hat{A}^{(L)}_{n})+L(1+\rho_{1})\beta^{L-1})$. Further, we have whp $\|\alpha^{(L)}\|_{2}\leq\beta^{L}$ (Lemma 5 in Appendix C), so that $w(\hat{A}^{(L)}_{n})\leq c\beta^{L}\sqrt{m}$. Then, for $\beta<1$, the bound $\beta^{L}\sqrt{m}+\beta^{L-1}L(1+\rho_{1})=O(1)$ if $L=\Omega(\frac{\log(\sqrt{m}+L(1+\rho_{1}))}{\log\frac{1}{\beta}})$. Interestingly, sufficient depth $L$ helps control the Gaussian width in the theoretical setting. ∎
Technicalities aside, Theorem 5 has two qualitative implications which are important to understand.
Remark 4.4.
Theorem 5 bounds the Gaussian width uniformly over all parameters in $B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$. Instead, we can consider the bound for parameters in $\Theta$ which are in the last 10% of the epochs of deep learning training. If the featurizer for such parameters $\theta\in\Theta$ are more stable, i.e., does not change much with changes in $W$, then the Gaussian width of the featurizer will be small. ∎
Remark 4.5.
The result implies that under suitable conditions, the Gaussian width of gradients over all $p=md+(L-1)m^{2}+m$ parameters can be reduced to the Gaussian width of the featurizer, which is $m$ dimensional. If the featurizer does feature selection, e.g., as in variants of invariant risk minimization (IRM) [5, 93] or sharpness aware minimization (SAM) [26, 4], then the Gaussian width of the featurizer will be small. On the other hand, if the featurizer has many spurious and/or random features [70, 93], then the Gaussian width will be large. ∎
@@ 4.2.1 Bounds for FeedForward Networks (FFNs)
Figure 2: $L_{0}$ norm and $L_{1}$ norm of featurizers for (a) ResNet18 on CIFAR-10, (b) FFN on CIFAR-10, and (c) CNN on Fashion-MNIST of the last 10% epochs close to convergence. We plot both the average and maximum norms among 5 repetitions for each experiment. The figures demonstrate that for these models, as approaching convergence, the featurizers has relatively low $L_{0}$ norm and $L_{1}$ norm, which further indicates a small LGGW. All experiments are run on A100 GPUs.

### v2

@@ 4 Gaussian Width of Deep Learning Gradients for Neural Networks
In this section, we initiate the study of loss gradient Gaussian width (LGGW) of deep learning gradients. As a warm up, we review that bounded $L_{2}$ norm, e.g., $\|\xi(\theta;z)\|_{2}\leq c$ for any $\theta\in\Theta$ and $z\in{\cal Z}$, with no additional structure, leads to a large Gaussian width since for such $\hat{\Xi}_{n}$, $w({\hat{\Xi}_{n}})=\mathbb{E}_{\mathbf{g}}[\sup_{\hat{\xi}_{n}}\langle\hat{\xi}_{n},\mathbf{g}\rangle]\leq c\mathbb{E}_{\mathbf{g}}[\langle\mathbf{g}/\|\mathbf{g}\|_{2},\mathbf{g}\rangle]=c\mathbb{E}_{\mathbf{g}}[\|\mathbf{g}\|_{2}]=O(\sqrt{p})$. We posit that the Gaussian width of deep learning gradients is much smaller and present a set of technical results in support. We show that under standard assumptions, single sample LGGW of such deep learning models can be bounded in terms of the Gaussian width of the featurizer, i.e., output of the last-but-one layer, for both feedforward and residual networks. The results emphasize the benefits of the stability of featurizers for deep learning.
4.1 Bounds Based on Empirical Geometry of Gradients
(a) Sorted gradient coordinate absolute values (linear scale).
(b) Sorted gradient coordinate absolute values (log scale).
(c) Sorted gradient coordinate absolute values (linear scale).
(d) Sorted gradient coordinate absolute values (log scale).
Figure 2: (a-b) Sorted gradient coordinate in linear scale (a) and log scale (b) for SGD over 100 epochs on CIFAR-10. Model: 4-hidden-layer ReLU network with 256 nodes on each layer, with around 1,055,000 parameters. (c-d) Sorted gradient coordinate in linear scale (c) and log scale (d) for SGD over 100 epochs on MNIST. Model: 2-layer ReLU with 128 nodes on each layer, with roughly 120,000 parameters. Y-axis is the absolute value of gradient coordinates, X-axis is sorted in increasing order.
In the context of deep learning, recent work has illustrated that empirically the gradients are considerably structured [34, 65, 66, 30, 48]: during optimization, the gradients are “simple” in the sense that the sorted absolute gradient components decay quickly and the components have large values only for a small subset of parameters (e.g., see Figure 2), and at convergence, almost all gradients are really close to zero, the so-called interpolation condition [14, 63, 56, 11]. Such observations have formed the basis of more memory or communication-efficient algorithms, e.g., based on top-$k$ count-sketch, truncated SGD, and heavy hitters based analysis [72, 40]. We augment these empirical observations with some examples of gradient sets which have small Gaussian widths by construction or by suitable composition properties.
Example 4.1 (Ellipsoid).
Consider $\hat{\Xi}_{n}=\{\hat{\xi}_{n}\in\mathbb{R}^{p}\mid\sum_{j=1}^{p}((\hat{\xi}_{n}[j])^{2}/a_{j}^{2}=\tilde{O}(1),a\in\mathbb{R}^{p}\}$. Then, the Gaussian width $w(\hat{\Xi}_{n})=\tilde{O}(\|\mathbf{a}\|_{2})$ [81]. Note that if $\|\mathbf{a}\|_{2}=\tilde{O}(1)$, then we have $w(\hat{\Xi}_{n})=\tilde{O}(1)$. For example, if the elements of $\mathbf{a}$ sorted in decreasing order satisfy $a_{(j)}\leq c_{3}/\sqrt{j}$, then $\|\mathbf{a}\|_{2}=\tilde{O}(1)$. ∎
Example 4.2 ($k$-support norm).
The $k$-support norm
$\|\xi\|_{k}^{sp}=\inf_{\sum_{i}u_{i}=\xi}\left\{\sum_{i}\|u_{i}\|_{2}\mid\|u_{i}\|_{0}\leq k\right\}$
is based on an infimum convolution over $k$-sparse vectors [96, 97, 18], with $k=1$ corresponding to the $L_{1}$ norm. Consider $\hat{\Xi}_{n}=\{\hat{\xi}_{n}\mid\|\hat{\xi}_{n}\|_{k}^{sp}\leq c_{0}\sqrt{k}\}$. Note that among many other things $\hat{\Xi}_{n}$ includes the infimum convolution of all $k$-sparse gradients $\hat{\xi}_{n}$ with each non-zero element bounded by $c_{0}$. By construction, $w(\hat{\Xi}_{n})\leq c_{1}k\sqrt{\log(p/k)}$ [18]. ∎
Example 4.3 (Union of sets).
If $\hat{\Xi}_{n}$ is an union of $p^{r}$ sets, i.e., polynomial in $p$, where each set has a $\tilde{O}(1)$ Gaussian width and each element has $\tilde{O}(1)$ $L_{2}$ norm, then $w(\hat{\Xi}_{n})=\tilde{O}(\sqrt{r})$ [98, 18]. Taking convex hull of such $\hat{\Xi}_{n}$ can only change its Gaussian width by a constant amount, i.e., $w(\text{conv}(\hat{\Xi}_{n}))\leq c_{2}w(\hat{\Xi}_{n})$ [81](Theorem 2.4.15). ∎
Example 4.4 (Minkowski sum of sets).
If $\hat{\Xi}_{n}$ is Minkowski sum of a union of $p^{r}$ sets as in Example 4.3 and a $\sqrt{k}$ radius $k$-support norm ball as in Example 4.2, then $w(\hat{\Xi}_{n})=\max(k,\sqrt{r})\tilde{O}(1)$. ∎
Based on how the gradients empirically look for deep networks (Figures  2), it is arguably fair to assume that the gradients live in $\hat{\Xi}_{n}$ with small Gaussian width. We prove explicit upper bounds of single-sample LGGW for FFNs and ResNets in Sections 4.2.1 and 4.2.2 respectively.
@@ 4.2 Gaussian Width Bounds for Neural Networks
Let $\theta_{0}:=(\operatorname{vec}(W_{0}^{(1)})^{\top},\ldots,\operatorname{vec}(W_{0}^{(L)})^{\top},\mathbf{v}_{0}^{\top})^{\top}$ denote the initial weights, and $w_{0,ij}^{(l)}$ denote $i,j$-th element of $W_{0}^{(l)}$ for $l\in[L]$. At initialization, $w_{0,ij}^{(l)}\sim{\cal N}(0,\sigma_{0}^{(l)}{}^{2})$ for $l\in[L]$ where $\sigma_{0}^{(1)}=\frac{\sigma_{1}}{2\left(1+\sqrt{\frac{\log m_{1}}{2m_{1}}}\right)}$, $\sigma_{0}^{(l)}=\frac{\sigma_{1}}{1+\sqrt{\frac{m_{l-1}}{m_{l}}}+\sqrt{\frac{2\log m_{l}}{m_{l}}}}$, $l\geq 2$, $\sigma_{1}>0$, and $\left\lVert\mathbf{v}_{0}\right\rVert_{2}=1$. For any input $\mathbf{x}\in{\cal X}\subseteq\mathbb{R}^{d}$, $\left\|\mathbf{x}\right\|_{2}=1$.
@@
Assumption 3 is satisfied by commonly used activation functions such as (non-smooth) ReLU [1] and (smooth) GeLU [39]. And Assumption 4 is regarding the random initialization of the weights, variants of which are typically used in most work in deep learning theory [41, 1, 38, 31, 80, 7, 23, 22].
@@
Figure 3: $L_{0}$ norm and $L_{1}$ norm of featurizers for (a) ResNet18 on CIFAR-10, (b) FFN on CIFAR-10, and (c) CNN on Fashion-MNIST of the last 10% epochs close to convergence. We plot both the average and maximum norms among 5 repetitions for each experiment. The figures demonstrate that for these models, as approaching convergence, the featurizers has relatively low $L_{0}$ norm and $L_{1}$ norm, which further indicates a small LGGW.
4.2.1 Bounds for FeedForward Networks (FFNs)
We consider $f$ to be an FFN given by
$f(\theta;x)=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m_{L}}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m_{1}}}W^{(1)}\mathbf{x}))))~{},$ (23)
where $W^{(1)}\in\mathbb{R}^{m_{1}\times d},W^{(l)}\in\mathbb{R}^{m_{l}\times m_{l-1}},l\in\{2,\ldots,L\}$ are the layer-wise weight matrices, $\mathbf{v}\in\mathbb{R}^{m_{L}}$ is the last layer vector, $\phi(\cdot)$ is the (pointwise) activation function, and the total set of parameters $\theta\in\mathbb{R}^{p},p=\sum_{k=1}^{L}m_{k}m_{k-1}+m_{L}$ is given by
in which $m_{0}=d$. Our analysis for both FFNs and ResNets will be for gradients over all parameters in a fixed radius spectral norm ball around the initialization $\theta_{0}$, which is more flexible than Frobenius norm balls typically used in the literature [86, 20, 95, 9]:
[…]
Let $B_{\rho}^{\operatorname{Spec}}(\theta_{0})\in\mathbb{R}^{p-m_{L}}$ be the corresponding set of weight matrices $W=\{W^{(l)},\ell\in[L]\}$. The main implications of our results and analyses go through as long as the layerwise spectral radius $\rho=o(\sqrt{m_{l}})$ for all $l\in[L]$, and last layer radius $\rho_{1}=O(1)$, which are arguably practical.
Let
$h^{(L)}(W,\mathbf{x})$ $:=\phi\left(\frac{1}{\sqrt{m_{L}}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m_{1}}}W^{(1)}\mathbf{x})))\right)~{}\in\mathbb{R}^{m_{L}}$ (26)
be the “featurizer,” i.e., output of the last but one layer, so that $f(\theta;\mathbf{x})=\mathbf{v}^{\top}h^{(L)}(W,\mathbf{x})$ with $\theta=(W,\mathbf{v})$. Consider the set of all featurizer values for a single sample:
$A^{(L)}(\mathbf{x})=\left\{h^{(L)}(W,\mathbf{x})\mid W\in B_{\rho}^{\operatorname{Spec}}(\theta_{0})\right\}$ (27)
We now establish an upper bound on single-sample LGGW of FFN gradient set $\Xi^{\operatorname{ffn}}$ corresponding to parameters in $B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$. In particular, we show that the single-sample LGGW can effectively be bounded by the Gaussian width of the featurizer set.
Theorem 4 (LGGW: FFNs).
Under Assumptions 3 and 4, with $\beta_{l}:=\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}$, $l\in[L]$, with probability at least $\left(1-\sum_{l=1}^{L}\frac{2}{m_{l}}\right)$ over the randomness of the initialization, we have
$w(\Xi^{\operatorname{ffn}})\leq c_{1}w(A^{(L)})+c_{2}(1+\rho_{1})\sqrt{m_{L}}\left(\prod_{l=1}^{L}\beta_{l}\right)\sum_{l=1}^{L}\frac{1}{\beta_{l}\sqrt{m_{l}}}~{}.$ (28)
Remark 4.1.
For the simplest case, with $m_{l}=m$ for all $l\in[L]$, we have that
$w(\Xi^{\operatorname{ffn}})\leq c_{1}w(A^{(L)})+c_{2}L(1+\rho_{1})\beta^{L-1},$
where $\beta=\sigma_{1}+\frac{\rho}{\sqrt{m}}$. Note that choosing $\sigma_{1}<1-\frac{\rho}{\sqrt{m}}$, i.e., mildly small initialization variance, satisfies $\beta<1$. When $\beta<1$, we have $w(\Xi^{\operatorname{ffn}})=O(1)$ as long as $L=\Omega(\frac{\log(\sqrt{m}+L(1+\rho_{1}))}{\log\frac{1}{\beta}})$, which holds for moderate depth $L$. To see this, note that we have $w(\Xi^{\operatorname{ffn}})=O(w(A^{(L)})+L(1+\rho_{1})\beta^{L-1})$. Further, we have whp $\|\alpha^{(L)}\|_{2}\leq\beta^{L}$ (Lemma 7 in Appendix C), so that $w(A^{(L)})\leq c\beta^{L}\sqrt{m}$. Then, for $\beta<1$, the bound $\beta^{L}\sqrt{m}+\beta^{L-1}L(1+\rho_{1})=O(1)$ if $L=\Omega(\frac{\log(\sqrt{m}+L(1+\rho_{1}))}{\log\frac{1}{\beta}})$. Interestingly, sufficient depth $L$ helps control the Gaussian width.∎
Remark 4.2.
In practice, the dimensions of the hidden layers $m_{l},l\in[L-1]$ are always much larger than the dimension of the output layer $m_{L}$, i.e., $m_{l}\gg m_{L}$, and our bound is even better under such situations. To see this, first note that $1-\sum_{l=1}^{L}\frac{2}{m_{l}}\gg 1-\frac{2L}{m_{L}}$. Second, for the second term in the bound which depends on parts other than the featurizer, $\sqrt{m_{L}}\prod_{l=1}^{L}\beta_{l}\sum_{l=1}^{L}\frac{1}{\beta_{l}\sqrt{m_{l}}}\ll L\beta_{L}^{L-1}$, since $\beta_{L}>\beta_{l}$. Therefore, the probability is higher and the bound is smaller when $m_{l}\gg m_{L}$, as in real neural networks. ∎
Remark 4.3.
The result implies that under suitable conditions, the Gaussian width of gradients over all $p=\sum_{k=1}^{d}m_{k}m_{k-1}+m_{L}$ parameters can be reduced to the Gaussian width of the featurizer, which is $m_{L}$ dimensional. If the featurizer does feature selection, e.g., as in variants of invariant risk minimization (IRM) [6, 94] or sharpness aware minimization (SAM) [26, 4], then the Gaussian width of the featurizer will be small. On the other hand, if the featurizer has many spurious and/or random features [71, 94], then the Gaussian width will be large. ∎